(3x-7)/(5x)=(1)/(4)
Answer by stanbon(75887)

(3x-7)/(5x)=(1)/(4)
----
Cross-multiply:
-----
4(3x-7) = 5x
---
12x - 28 = 5x
7x = 28
x = 4
---------------
Cheers,
Stan H.
===============